OpenStudy (mackenzie2013):

Which of the following is equal to 5 4/9? 5.49 5.25 7.25 5.4

OpenStudy (mackenzie2013):

Got it nevermind ;)

OpenStudy (compassionate):

Hi, 5 * 4/9 These are simple, just multiply your numerator by 5. So 4 * 5/9 = 20/9 = 2.22